OpenStudy (anonymous):

Factor completely: 4x2 + 25x + 6 (4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1) 3. Factor completely: 3x2 + 2x − 1 (3x + 1)(x − 1) (3x + 1)(x + 1) (3x − 1)(x + 1) (3x − 1)(x − 1) 4. Factor completely: 18x2 − 21x − 15 3(2x + 1)(3x − 5) 3(2x − 5)(3x + 1) 3(2x − 1)(3x + 5) 3(6x + 1)(x − 5) 5. Factor completely: 3x2 + 5x + 1 (3x + 1)(x + 1) (3x + 5)(x + 1) (3x − 5)(x + 1) Prime

OpenStudy (anonymous):

$$4x^2+25x+6=4x^2+24x+x+6=4x(x+6)+1(x+6)=(4x+1)(x+6)$$

OpenStudy (anonymous):

Can you do something similar for the rest?

OpenStudy (mertsj):

These look like test questions.

OpenStudy (anonymous):

they arent they are study questions that the teacher gave me and i dont even know how to do

OpenStudy (mertsj):

Follow the example shown above.

OpenStudy (anonymous):

i dont even know how to do it lol i havent learned it

OpenStudy (mertsj):

Then explain that to your teacher and ask him/her to go over it again.